OpenStudy (anonymous):

What is the quotient 6-x/x^2 + 2x - 3 divided by x^2 - 4x - 12/x^2+4x+3 in simplified form? State any restrictions on the variable

OpenStudy (anonymous):

I know the answer already is -(x+1)/(x-1)(x+2)

OpenStudy (anonymous):

how do I find the variables?

OpenStudy (luigi0210):

Do you know how to start off?

OpenStudy (ashleyisakitty):

@Luigi0210 how do i do this one? :P

OpenStudy (luigi0210):

change the division to multiplication and flip the second

OpenStudy (luigi0210):

Do you need me to do it to show you?

OpenStudy (ashleyisakitty):

Yes, that would be most helpful to go by steps :)

OpenStudy (luigi0210):

Okay so we have this: \[\frac{6-x}{x^2+2x-3}\div\frac{x^2-4x-12}{x^2+4x+3}\] So now you change the signs and flip the equation around: \[\frac{6-x}{x^2+2x-3}*\frac{x^2+4x+3}{x^2-4x-12}\] So now start factoring what you can: \[\frac{-(x-6)}{(x+3)(x-1)}*\frac{(x+3)(x+1)}{(x-6)(x+2)}\] Do you know what to do now Miss Ashley?

OpenStudy (ashleyisakitty):

yes! thank you!

OpenStudy (luigi0210):

Anytime :)